# Towards a generic absoluteness theorem for Chang models

### v2

@@ 5.3. Making universally Baire sets generic: The One Step Construction
    a $\lambda$-ground $U$ of $V$ such that $U\models``\eta$ is uB-preserving”, and
@@
    $W\in U_{\chi}$ such that in $U$, $W$ is $(\chi,\eta,\lambda)$-saturated and infinitely tamed above $\eta$ [^25] See Definition 4.9 and Definition 3.1..
Suppose ${\sf a}$ is a $(U,\chi,\lambda,\{\eta,W\})$-realizable rock. In this section, we will use supercript ${\sf a}$ to denote the preimages of objects in the range of $\sigma^{\sf a}$. Thus, $W^{\sf a}$ is the $\sigma^{\sf a}$-preimage of $W$. We will also use this notation with elementary embeddings as well. Thus, if $r:N\rightarrow U_{\chi}$ then $W^{N}$ is the $r$-preimage of $W$ (if it exists). The notation depends on the embedding as well, but omitting this dependence will not cause confusion.
@@
We once again remark that the correct terminology should be that ${\sf{b}}$ b-catches $A$ relative to $(U,\chi,\eta,\lambda,W)$, but since this object is fixed in this section we will omit it from our terminology. In order to show that universally Baire sets end up in derived models of rocks, we need to develop a method for b-catching universally Baire sets. The next lemma shows how to convert t-catching to b-catching. It is not too hard to t-catch universally Baire sets but this must be done in a way that preserves the fact that other sets have been b-caught. This is the step that requires the UB-Capturing Principle, and it will appear in applications of Lemma 5.16. The reader may profit from reviewing Definition 3.4.
@@
(2) is a consequence of the standard copying construction. Indeed, pick $\alpha<\operatorname{lh}(E)$ and let $id$ be the identity function on $\kappa$. We then have that $\alpha=\pi_{E}^{N}(id)(a)$. Hence,
@@
Because ${\rm cp}(\sigma^{{\sf{b}},{\sf d}})\geq\kappa^{\sf{b}}$ and because $\sigma^{\sf{b}}\,"\bar{\mu}=\sigma^{\sf d}\,"(\sigma^{{\sf{b}},{\sf d}}\,"\bar{\mu})$, we only need to prove that $\sigma^{{\sf{b}},{\sf d}}\,"\bar{\mu}\in M^{\sf d}[k]$. To see this, notice that $\sigma^{{\sf{b}},{\sf d}}$ can be lifted to $M^{\sf{b}}[k]$ (because ${\rm cp}(\sigma^{{\sf{b}},{\sf d}})\geq\kappa^{\sf{b}}$ and $\beta<\kappa^{\sf{b}}$). Let $j:M^{\sf{b}}[k]\rightarrow M^{\sf d}[k]$ be this embedding. We thus have that $j\restriction M^{\sf{b}}=\sigma^{{\sf{b}},{\sf d}}$. It then follows that $\sigma^{\sf{b}}\,"\bar{\mu}=j(\bar{\mu})=j\,"\bar{\mu}$ (because $\bar{\mu}$ is a countable set). ∎
@@ 5.4. Summary
    $U$ be a $\lambda$-ground of $V$ such that $U\models``\eta$ is uB-preserving”,
@@
Clearly $A$ may not have a $\sigma^{\sf a}$-code. Next we introduce $A$-genericity pairs.
@@
    ${\rm cp}(\sigma^{{\sf a},{\sf{b}}})>\delta_{1}^{\sf a}$,
@@
Suppose $A$ is a universally Baire set, $(y,z)$ is a code of $A$ and ${\sf{b}}$ is an ${\sf a}$-realizable rock. Then ${\sf{b}}$ t-captures $A$ (relative to $(U,\chi,\eta,\lambda,W)$).
@@
Let $(p,(\delta_{i}:i\leq 3))$ witness that $(y,z)$ is an ${\sf a}$-code and $k\subseteq\operatorname{Col}(\omega,\delta_{2}^{\sf{b}})$ be $M^{\sf{b}}$-generic such that $(y,z)\in M^{\sf{b}}$. Let then $\bar{\nu}=(u_{y}(i):i\in z)$. We have that $\bar{\nu}\in M^{\sf{b}}[k]$ and $A=S_{\sigma^{\sf{b}}\,"\bar{\nu}}$. Let $v=(f_{1}\circ f_{0})^{-1}$ and set $\bar{\mu}=v^{{\sf{b}}}(\bar{t})$. Then $\bar{\mu}\in M^{\sf{b}}[k]$ and $\sigma^{\sf{b}}\,"\bar{\mu}\subseteq W_{2}(=W)$. To show that $(k,\bar{\mu})$ witness that ${\sf{b}}$ t-catches $A$, it is enough to show that $A=S_{\sigma^{\sf{b}}\,"\bar{\mu}}$. This, however, easily follows from the fact that $(f_{1},W_{2},W_{1})$ and $(f_{0},W_{1},W_{0})$ are flipping systems. ∎
If $({\sf{b}},E)$ is a $(A\mid y,z)$-genericity ${\sf a}$-realization then we will informally say that $E$ is the $A$-sealing extender or $E$ seals $A$. The reason for the terminology is that we will show that $A$ is in the derived model of $M^{\sf a}_{E}$. We will obtain an $A$-genericity ${\sf a}$-realization $({\sf{b}},E)$ by performing genericity iterations. In particular, ${\sf{b}}$ will be an iterate of ${\sf a}$ via an iteration tree ${\mathcal{T}}$ of length $\omega+1$. The following theorem is the summary of what we have shown. For convenience, we will re-introduce all previously fixed objects. The reader may wish to review Definition 5.14.
@@
Suppose $\chi>\eta>\lambda$ are inaccessible cardinals, the UB-Capturing Principle holds at $(\chi,\eta,\lambda)$ as witnessed by $(X,W)$. Let $W^{\prime}\in X$ be as in Clause 3 of Definition 4.7 and $g$ be a $<\lambda$-generic such that ${|X|}^{V[g]}=\aleph_{0}$. Then the following holds in $V[g]$ [^33] We apply the concepts we have developed to $U=V$ and $V=V[g]$. Clearly, $U$ is a $\lambda$-ground of $V$..
Suppose $B\in\Gamma^{\infty}_{g}$, $\vec{A}\in(\Gamma^{\infty}_{g})^{<\omega_{1}}$ [^34] This is $\omega_{1}$ of $V[g]$., $z\subseteq W$ is countable, $a\subseteq\eta$ is bounded and ${\sf a}$ is the rock induced by $X(z\cup a)$. Assume ${\sf a}$ b-catches every member of $\vec{A}$ (relative to $(U,\chi,\eta,\lambda,W^{\prime})$). There is then an ${\sf a}$-realizable ${\sf{b}}$, an ${\sf a}$-realizable ${\sf e}$ and $E\in M^{\sf{b}}$ such that
@@
    ${\sf{b}}$ t-catches $B$ (relative to $(U,\chi,\eta,\lambda,W^{\prime})$),
@@
    ${\sf e}$ b-catches both $A$ and $B$ (relative to $(U,\chi,\eta,\lambda,W^{\prime})$).
@@
The proof is an application of Lemmas 5.15 and 5.16. First, applying Definition 4.7, we find a countable $z^{\prime}\subseteq W$ such that $(X(z\cup z^{\prime}\cup a),W)$ captures $B$. Let ${\sf{b}}^{\prime}$ be the rock induced by $X(z\cup z^{\prime}\cup a)$. Because $X$ is $(W\mid\eta)$-full, we have that ${\sf{b}}^{\prime}$ is ${\sf a}$-realizable and $\sigma^{{\sf a},{\sf{b}}^{\prime}}\restriction(M^{\sf a}\cap{\wp}(\kappa^{\sf a}))=id$.
@@
    $p=(W_{2},W_{1},W_{0},f_{1},f_{0})\in V_{\chi}$ is a supporting system that flips through $(\delta_{i}:i\leq 3)$ [^35] See Definition 5.11. and
@@
Next, using Neeman’s Genericty Theorem, we find a ${\sf{b}}^{\prime}$-realizable ${\sf{b}}$ such that ${\rm cp}(\sigma^{{\sf{b}}^{\prime},{\sf{b}}})>\delta^{{\sf{b}}^{\prime}}_{1}$ and for some $M^{{\sf{b}}}$-generic $k\subseteq\operatorname{Col}(\omega,\delta_{2}^{{\sf{b}}})$, $(r,s)\in M^{{\sf{b}}}[k]$. It now follows from Lemma 5.20 that ${\sf{b}}$ t-catches $B$.
At this stage, we have that ${\sf{b}}$ is ${\sf a}$-realizable, t-catches $B$ and $\sigma^{{\sf a},{\sf{b}}}\restriction(M^{\sf a}\cap{\wp}(\kappa^{{\sf a}}))=id$. Applying Lemma 5.15 we get an extender $E\in M^{\sf{b}}$ such that ${\rm cp}(E)=\kappa^{{\sf a}}$, $W_{2}^{{\sf{b}}}\in M^{\sf{b}}|\operatorname{lh}(E)$ and $M^{\sf{b}}\models``E$ is a nice wf-preserving for $W_{2}^{{\sf{b}}}$ extender"""" and an ${\sf a}$-realizable ${\sf e}$ such that $M^{\sf e}=Ult({\mathcal{M}}^{\sf a},E)$, $\sigma^{{\sf a},{\sf e}}=\pi_{E}^{M^{\sf a}}$ and ${\sf e}$ b-catches $B$. Finally it follows from Lemma 5.16 that ${\sf e}$ b-catches every member of $\vec{A}$. ∎
@@
Suppose $\chi>\eta>\lambda$ are inaccessible cardinals, $U$ is a $\lambda$-ground and $U\models``$the UB-Capturing Principle holds at $(\chi,\eta,\lambda)$ as witnessed by $(X,W)$”. Suppose $i\in\{0,1\}$, $k\subseteq\operatorname{Col}(\omega_{i},{|\Gamma^{\infty}|})$ is $V$-generic and $W^{\prime}\in X$ is such that $W\subseteq W^{\prime}$ and $W^{\prime}$ witnesses Clause 3 of Definition 4.7 applied to $(X,W)$. Suppose $\iota\leq\omega_{1}$ is a limit ordinal and $k^{\prime}\in V[k]$ is a surjection $k^{\prime}:\iota\rightarrow\Gamma^{\infty}$ with the property that for all $\alpha<\iota$, $k^{\prime}\restriction\alpha\in V$. For $i\in\{0,1\}$, we say $D=(({\sf a}_{\alpha},{\sf{b}}_{\alpha},E_{\alpha})\mid\alpha<\iota)$ is a $\sf{DM}$-sequence for $V$ supported by $(X,W,W^{\prime},U,\chi,\eta,\lambda,k^{\prime})$ if in $V[k]$,
@@
    for $\alpha<\iota$, ${\sf a}_{\alpha}$ b-catches every member of $k\restriction\alpha$ relative to $(U,\chi,\eta,\lambda,W^{\prime})$,
@@
    for every $\alpha<\iota$, $({\sf{b}}_{\alpha},E_{\alpha},{\sf a}_{\alpha+1})$ is a possible output of the one-step-construction applied to $({\sf a}_{\alpha},k^{\prime}\restriction\alpha,k^{\prime}(\alpha),X,W,W^{\prime},U,\chi,\eta,\lambda)$ [^37] Thus, ${\sf{b}}_{\alpha}$ t-catches $A_{\alpha}$, $E$ is wf-preserving extender for some $W_{2}^{\sf{b}}$ as in the proof of Theorem 5.21 and ${\sf a}_{\alpha+1}$ b-catches every member of $k\restriction\alpha+1$.,
@@
    for every $A\in\Gamma^{\infty}$ [^38] While this statement is evaluated in $V[g]$, it applies to $A\in V$. there is an $\alpha<\iota$ such that if ${\sf a}_{\alpha}=(\sigma_{X_{\alpha}},M_{X_{\alpha}},\kappa_{X_{\alpha}})$ is the rock induced by $X_{\alpha}$ then ${\sf a}_{\alpha}$ b-catches $A$ relative to $(U,\chi,\eta,\lambda,W^{\prime})$ [^39] Because if $x\in\mathbb{R}$ then $\{x\}$ is a universally Baire set, we have that for each $x\in\mathbb{R}$ there is an $\alpha<\iota$ and an $M^{{\sf a}_{\alpha}}$-generic $g\subseteq\operatorname{Col}(\omega,{<}\kappa^{{\sf a}_{\alpha}})$ such that $x\in M^{{\sf a}_{\alpha}}[g]$.,
@@
    for every limit ordinal $\alpha<\iota$, ${\sf a}_{\alpha}$ is a direct limit of $({\sf a}_{\beta}\mid\beta<\alpha)$ [^40] More precisely, ${\sf a}_{\alpha}$ is the ${\sf a}_{0}$-realizable rock induced by $\cup_{\beta<\alpha}\operatorname{rng}(\sigma^{\sf a}_{\alpha})$..
@@
The following is our main theorem on the existence of $\sf{DM}$-sequences. The reader may wish to review Definition 2.8, Definition 3.4, Lemma 3.5, Definition 4.7 (The UB-Capturing Principle), Theorem 4.10, Definition 5.11, Definition 5.18, Lemma 5.15 and Theorem 5.21. We state two versions of the theorem. In the first version, we collapse ${|\Gamma^{\infty}|}$ to be $\omega$ and in the second, we collapse it to be $\omega_{1}$. The proofs are straightforward application of Theorem 5.21 and Lemma 5.16, and we leave them to the reader.
@@
Suppose $\chi>\eta>\lambda$ are inaccessible cardinals, $U$ is a $\lambda$-ground, $U\models``$the UB-Capturing Principle holds at $(\chi,\eta,\lambda)$ as witnessed by $(X,W)$”, and $W^{\prime}\in X$ is such that $W\subseteq W^{\prime}$ and $W^{\prime}$ witnesses Clause 3 of Definition 4.7 applied to $(X,W)$. Suppose $k\subseteq\operatorname{Col}(\omega,{|\Gamma^{\infty}|})$ is $V$-generic, $\iota<\omega_{1}$ is a limit ordinal and $k^{\prime}:\iota\rightarrow\Gamma^{\infty}$ is a surjection in $V[k]$ with the property that for all $\alpha<\iota$, $k^{\prime}\restriction\alpha\in V$. Then in $V[k]$, there is a $\sf{DM}$-sequence $D=(({\sf a}_{\alpha},{\sf{b}}_{\alpha},E_{\alpha})\mid\alpha<\iota)$ for $V$ supported by $(X,W,W^{\prime},U,\chi,\eta,\lambda,k^{\prime})$.
@@
Suppose $\chi>\eta>\lambda$ are inaccessible cardinals, $U$ is a $\lambda$-ground, $U\models``$the UB-Capturing Principle holds at $(\chi,\eta,\lambda)$ as witnessed by $(X,W)$” and $W^{\prime}\in X$ is such that $W\subseteq W^{\prime}$ and $W^{\prime}$ witnesses Clause 3 of Definition 4.7 applied to $(X,W)$. Suppose $k\subseteq\operatorname{Col}(\omega_{1},{|\Gamma^{\infty}|})$ is $V$-generic. Then in $V[k]$, there is a $\sf{DM}$-sequence $D=(({\sf a}_{\alpha},{\sf{b}}_{\alpha},E_{\alpha})\mid\alpha<\iota)$ for $V$ supported by $(X,W,W^{\prime},U,\chi,\eta,\lambda,k^{\prime})$.
@@
is the underlying sequence of ${\sf{DM}}$ if for $\alpha<\beta<\iota$, $M_{\alpha}=M^{{\sf a}_{\alpha}}$, $\sigma_{\alpha,\beta}=\sigma^{{\sf a}_{\alpha},{\sf a}_{\beta}}$, $\sigma_{\alpha}=\sigma^{{\sf a}_{\alpha}}$, $\kappa_{\alpha}=\kappa^{{\sf a}_{\alpha}}$, $M_{\iota}$ is the direct limit of $(M_{\alpha},\sigma_{\alpha,\beta}:\alpha<\beta<\iota)$, $\sigma_{\alpha,\iota}:M_{\alpha}\rightarrow M_{\iota}$ is the direct limit embedding, and $\sigma_{\iota}:M_{\iota}\rightarrow U_{\chi}$ is the canonical realizable embedding given by $\sigma_{\iota}(x)=\sigma_{\alpha}(\bar{x})$ where $\alpha<\iota$ and $\bar{x}\in M_{\alpha}$ are such that $\sigma_{\alpha,\iota}(\bar{x})=x$.

### v3

@@ 5.3. Making universally Baire sets generic: The One Step Construction
    a $\lambda$-ground $U$ of $V$ such that $U\models``\eta$ is uB-preserving”, [^24] See Definition 3.5. and
@@
    $W\in U_{\chi}$ such that in $U$, $W$ is $(\chi,\eta,\lambda)$-saturated and infinitely tamed above $\eta$ [^25] See Definition 4.9 and Definition 3.1.
Suppose ${\sf a}$ is a $(U,\chi,\lambda,\{\eta,W\})$-realizable rock. In this section, we will use superscript ${\sf a}$ to denote the preimages of objects in the range of $\sigma^{\sf a}$. Thus, $W^{\sf a}$ is the $\sigma^{\sf a}$-preimage of $W$. We will use this notation with elementary embeddings as well. Thus, if $r:N\rightarrow U_{\chi}$ then $W^{N}$ is the $r$-preimage of $W$ (if it exists). The notation depends on the embedding as well, but omitting this dependence will not cause confusion.
@@
We once again remark that the correct terminology should be that ${\sf{b}}$ b-catches $A$ relative to $(U,\chi,\lambda,\{\eta,W\})$, but since this object is fixed in this section we will omit it from our terminology. In order to show that universally Baire sets end up in derived models of rocks, we need to develop a method for b-catching universally Baire sets. The next lemma shows how to convert t-catching to b-catching. It is not too hard to t-catch universally Baire sets but this must be done in a way that preserves the fact that other sets have been b-caught. This is the step that requires the UB-Capturing Principle, and it will appear in applications of Lemma 5.16. The reader may profit from reviewing Definition 3.5.
@@
(2) is a consequence of the standard copying construction. Indeed, pick $\alpha<\operatorname{lh}(E)$ and let $id$ be the identity function on $\kappa$. We then have that $\alpha=\pi_{E}^{N}(id)(\alpha)$. Hence,
@@
Because ${\rm cp}(\sigma^{{\sf{b}},{\sf d}})\geq\kappa^{\sf{b}}$ and because $\sigma^{\sf{b}}\,"\bar{\mu}=\sigma^{\sf d}\,"(\sigma^{{\sf{b}},{\sf d}}\,"\bar{\mu})$, we only need to prove that $\sigma^{{\sf{b}},{\sf d}}\,"\bar{\mu}\in M^{\sf d}[k]$. To see this, notice that $\sigma^{{\sf{b}},{\sf d}}$ can be lifted to $M^{\sf{b}}[k]$ (because ${\rm cp}(\sigma^{{\sf{b}},{\sf d}})\geq\kappa^{\sf{b}}$ and $\beta<\kappa^{\sf{b}}$). Let $j:M^{\sf{b}}[k]\rightarrow M^{\sf d}[k]$ be this embedding. We thus have that $j\restriction M^{\sf{b}}=\sigma^{{\sf{b}},{\sf d}}$. It then follows that $\sigma^{{\sf{b}},{\sf d}}\,"\bar{\mu}=j\,"\bar{\mu}=j(\bar{\mu})\in M^{\sf d}[k]$ (because $\bar{\mu}$ is a countable set in $M^{\sf{b}}[k]$). ∎
@@ 5.4. Summary
    $U$ be a $\lambda$-ground of $V$ such that $U\models``\eta$ is uB-preserving”, [^31] See Definition 3.5.
@@
Clearly $A$ may not have an ${\sf a}$-code. Next we introduce $A$-genericity pairs.
@@
    ${\rm cp}(\sigma^{{\sf a},{\sf{b}}})>\delta_{1}^{\sf a}$, [^34] Recall that $\delta_{1}^{\sf a}=(\sigma^{\sf a})^{-1}(\delta_{1})$, etc.
@@
Suppose $A$ is a universally Baire set, $(y,z)$ is a code of $A$ and $({\sf{b}},E)$ is an $(A\mid y,z)$-genericity ${\sf a}$-realization. Then ${\sf{b}}$ t-catches $A$ (relative to $(U,\chi,\lambda,\{\eta,W\})$).
@@
Let $(p,(\delta_{i}:i\leq 3))$ witness that $(y,z)$ is an ${\sf a}$-code and $k\subseteq\operatorname{Col}(\omega,\delta_{2}^{\sf{b}})$ be $M^{\sf{b}}$-generic such that $(y,z)\in M^{\sf{b}}[k]$. Let then $\bar{\nu}=(u_{y}(i):i\in z)$ be as in Definition 5.18. We have that $\bar{\nu}\in M^{\sf{b}}[k]$ and $A=S_{\sigma^{\sf{b}}\,"\bar{\nu}}$. Let $v=(f_{1}\circ f_{0})^{-1}$ and set $\bar{\mu}=v^{{\sf{b}}}(\bar{\nu})$. Then $\bar{\mu}\in M^{\sf{b}}[k]$ and $\sigma^{\sf{b}}\,"\bar{\mu}\subseteq W_{2}(=W)$. To show that $(k,\bar{\mu})$ witnesses that ${\sf{b}}$ t-catches $A$, it is enough to show that $A=S_{\sigma^{\sf{b}}\,"\bar{\mu}}$. This, however, easily follows from the fact that $(f_{1},W_{2},W_{1})$ and $(f_{0},W_{1},W_{0})$ are flipping systems. ∎
If $({\sf{b}},E)$ is an $(A\mid y,z)$-genericity ${\sf a}$-realization then we will informally say that $E$ is the $A$-sealing extender or $E$ seals $A$. The reason for the terminology is that we will show that $A$ is in the derived model of $M^{\sf a}_{E}$. We will obtain an $A$-genericity ${\sf a}$-realization $({\sf{b}},E)$ by performing genericity iterations. In particular, ${\sf{b}}$ will be an iterate of ${\sf a}$ via an iteration tree ${\mathcal{T}}$ of length $\omega+1$. The following theorem is the summary of what we have shown. For convenience, we will re-introduce all previously fixed objects. The reader may wish to review Definition 5.14.
@@
Suppose $\chi>\eta>\lambda$ are inaccessible cardinals and the UB-Capturing Principle holds at $(\chi,\eta,\lambda)$ as witnessed by $(X,W)$. Let $W^{\prime}\in X$ be as in Clause (3) of Definition 4.7 and $g$ be ${<}\lambda$-generic such that ${|X|}^{V[g]}=\aleph_{0}$. Then the following holds in $V[g]$ [^35] We apply the concepts we have developed to $U=V$ and $V=V[g]$. Clearly, $U$ is a $\lambda$-ground of $V$..
Suppose $B\in\Gamma^{\infty}_{g}$, $\vec{A}\in(\Gamma^{\infty}_{g})^{<\omega_{1}}$ [^36] This is $\omega_{1}$ of $V[g]$., $z\subseteq W$ is countable, $a\subseteq\eta$ is bounded and ${\sf a}$ is the rock induced by $X(z\cup a)$. Assume ${\sf a}$ b-catches every member of $\vec{A}$ (relative to $(U,\chi,\lambda,\{\eta,W^{\prime}\})$). There is then an ${\sf a}$-realizable ${\sf{b}}$, an ${\sf a}$-realizable ${\sf e}$ and $E\in M^{\sf{b}}$ such that
@@
    ${\sf{b}}$ t-catches $B$ (relative to $(U,\chi,\lambda,\{\eta,W^{\prime}\})$),
@@
    ${\sf e}$ b-catches every member of $\vec{A}$ and $B$ (relative to $(U,\chi,\lambda,\{\eta,W^{\prime}\})$).
@@
The proof is an application of Lemmas 5.15 and 5.16. First, applying Definition 4.7, we find a countable $z^{\prime}\subseteq W$ such that $(X(z\cup z^{\prime}\cup a),W)$ captures $B$. Let ${\sf{b}}^{\prime}$ be the rock induced by $X(z\cup z^{\prime}\cup a)$. Because $X$ is $(W\mid\eta)$-full, [^37] See Definition 4.5. we have that ${\sf{b}}^{\prime}$ is ${\sf a}$-realizable and $\sigma^{{\sf a},{\sf{b}}^{\prime}}\restriction(M^{\sf a}\cap{\wp}(\kappa^{\sf a}))=id$.
@@
    $p=(W_{2},W_{1},W_{0},f_{1},f_{0})\in V_{\chi}$ is a supporting system that flips through $(\delta_{i}:i\leq 3)$, [^38] See Definition 5.11. and
@@
Next, using Neeman’s Genericity Theorem, we find a ${\sf{b}}^{\prime}$-realizable ${\sf{b}}$ such that ${\rm cp}(\sigma^{{\sf{b}}^{\prime},{\sf{b}}})>\delta^{{\sf{b}}^{\prime}}_{1}$ and for some $M^{{\sf{b}}}$-generic $k\subseteq\operatorname{Col}(\omega,\delta_{2}^{{\sf{b}}})$, $(r,s)\in M^{{\sf{b}}}[k]$. It now follows from Lemma 5.20 that ${\sf{b}}$ t-catches $B$.
At this stage, we have that ${\sf{b}}$ is ${\sf a}$-realizable, t-catches $B$ and $\sigma^{{\sf a},{\sf{b}}}\restriction(M^{\sf a}\cap{\wp}(\kappa^{{\sf a}}))=id$. Applying Lemma 5.15 we get an extender $E\in M^{\sf{b}}$ such that ${\rm cp}(E)=\kappa^{{\sf a}}$, $W_{2}^{{\sf{b}}}\in M^{\sf{b}}|\operatorname{lh}(E)$ and $M^{\sf{b}}\models``E$ is a nice wf-preserving for $W_{2}^{{\sf{b}}}$ extender"""" and an ${\sf a}$-realizable ${\sf e}$ such that $M^{\sf e}=Ult(M^{\sf a},E)$, $\sigma^{{\sf a},{\sf e}}=\pi_{E}^{M^{\sf a}}$ and ${\sf e}$ b-catches $B$. Finally it follows from Lemma 5.16 that ${\sf e}$ b-catches every member of $\vec{A}$. ∎
@@
Suppose $\chi>\eta>\lambda$ are inaccessible cardinals, $U$ is a $\lambda$-ground and $U\models``$the UB-Capturing Principle holds at $(\chi,\eta,\lambda)$ as witnessed by $(X,W)$”. Suppose $i\in\{0,1\}$, $k\subseteq\operatorname{Col}(\omega_{i},{|\Gamma^{\infty}|})$ is $V$-generic and $W^{\prime}\in X$ is such that $W\subseteq W^{\prime}$ and $W^{\prime}$ witnesses Clause (3) of Definition 4.7 applied to $(X,W)$. Suppose $\iota\leq\omega_{1}$ is a limit ordinal and $k^{\prime}\in V[k]$ is a surjection $k^{\prime}:\iota\rightarrow\Gamma^{\infty}$ with the property that for all $\alpha<\iota$, $k^{\prime}\restriction\alpha\in V$. We say $D=(({\sf a}_{\alpha},{\sf{b}}_{\alpha},E_{\alpha})\mid\alpha<\iota)$ is a $\sf{DM}$-sequence for $V$ supported by $(X,W,W^{\prime},U,\chi,\eta,\lambda,k^{\prime})$ if in $V[k]$,
@@
    for $\alpha<\iota$, ${\sf a}_{\alpha}$ b-catches every member of $k\restriction\alpha$ relative to $(U,\chi,\lambda,\{\eta,W^{\prime}\})$,
@@
    for every $\alpha<\iota$, $({\sf{b}}_{\alpha},E_{\alpha},{\sf a}_{\alpha+1})$ is a possible output of the one-step-construction applied to $({\sf a}_{\alpha},k^{\prime}\restriction\alpha,k^{\prime}(\alpha),X,W,W^{\prime},U,\chi,\eta,\lambda)$, [^40] Thus, ${\sf{b}}_{\alpha}$ t-catches $A_{\alpha}$, $E_{\alpha}$ is a wf-preserving extender for some $W_{2}^{\sf{b}}$ as in the proof of Theorem 5.21 and ${\sf a}_{\alpha+1}$ b-catches every member of $k\restriction\alpha+1$.
@@
    for every $A\in\Gamma^{\infty}$ [^41] While this statement is evaluated in $V[g]$, it applies to $A\in V$. there is an $\alpha<\iota$ such that if ${\sf a}_{\alpha}=(\sigma_{X_{\alpha}},M_{X_{\alpha}},\kappa_{X_{\alpha}})$ is the rock induced by $X_{\alpha}$ then ${\sf a}_{\alpha}$ b-catches $A$ relative to $(U,\chi,\lambda,\{\eta,W^{\prime}\})$, [^42] Since $\{x\}$ is a universally Baire set for every $x\in\mathbb{R}$, we have that for each $x\in\mathbb{R}$ there is an $\alpha<\iota$ and an $M^{{\sf a}_{\alpha}}$-generic $g\subseteq\operatorname{Col}(\omega,{<}\kappa^{{\sf a}_{\alpha}})$ such that $x\in M^{{\sf a}_{\alpha}}[g]$.
@@
    for every limit ordinal $\alpha<\iota$, ${\sf a}_{\alpha}$ is a direct limit of $({\sf a}_{\beta}\mid\beta<\alpha)$. [^43] More precisely, ${\sf a}_{\alpha}$ is the ${\sf a}_{0}$-realizable rock induced by $\cup_{\beta<\alpha}\operatorname{rng}(\sigma^{\sf a}_{\alpha})$.
@@
The following is our main theorem on the existence of $\sf{DM}$-sequences. The reader may wish to review Definition 2.8, Definition 3.5, Lemma 3.6, Definition 4.7 (The UB-Capturing Principle), Theorem 4.10, Definition 5.11, Definition 5.18, Lemma 5.15 and Theorem 5.21. We state two versions of the theorem. In the first version, we collapse ${|\Gamma^{\infty}|}$ to be $\omega$ and in the second, we collapse it to be $\omega_{1}$. The proofs are straightforward applications of Theorem 5.21 and Lemma 5.16, and we leave them to the reader.
@@
Suppose $\chi>\eta>\lambda$ are inaccessible cardinals, $U$ is a $\lambda$-ground, $U\models``$the UB-Capturing Principle holds at $(\chi,\eta,\lambda)$ as witnessed by $(X,W)$”, and $W^{\prime}\in X$ is such that $W\subseteq W^{\prime}$ and $W^{\prime}$ witnesses Clause (3) of Definition 4.7 applied to $(X,W)$. Suppose $k\subseteq\operatorname{Col}(\omega,{|\Gamma^{\infty}|})$ is $V$-generic, $\iota<\omega_{1}$ is a limit ordinal and $k^{\prime}:\iota\rightarrow\Gamma^{\infty}$ is a surjection in $V[k]$ with the property that for all $\alpha<\iota$, $k^{\prime}\restriction\alpha\in V$. Then, in $V[k]$, there is a $\sf{DM}$-sequence $D=(({\sf a}_{\alpha},{\sf{b}}_{\alpha},E_{\alpha})\mid\alpha<\iota)$ for $V$ supported by $(X,W,W^{\prime},U,\chi,\eta,\lambda,k^{\prime})$.
@@
Suppose $\chi>\eta>\lambda$ are inaccessible cardinals, $U$ is a $\lambda$-ground, $U\models``$the UB-Capturing Principle holds at $(\chi,\eta,\lambda)$ as witnessed by $(X,W)$” and $W^{\prime}\in X$ is such that $W\subseteq W^{\prime}$ and $W^{\prime}$ witnesses Clause (3) of Definition 4.7 applied to $(X,W)$. Suppose $k\subseteq\operatorname{Col}(\omega_{1},{|\Gamma^{\infty}|})$ is $V$-generic. Then, in $V[k]$, there is a $\sf{DM}$-sequence $D=(({\sf a}_{\alpha},{\sf{b}}_{\alpha},E_{\alpha})\mid\alpha<\omega_{1})$ for $V$ supported by $(X,W,W^{\prime},U,\chi,\eta,\lambda,k)$.
@@
is the underlying sequence of $D$ if for $\alpha<\beta<\iota$, $M_{\alpha}=M^{{\sf a}_{\alpha}}$, $\sigma_{\alpha,\beta}=\sigma^{{\sf a}_{\alpha},{\sf a}_{\beta}}$, $\sigma_{\alpha}=\sigma^{{\sf a}_{\alpha}}$, $\kappa_{\alpha}=\kappa^{{\sf a}_{\alpha}}$. Moreover, $M_{\iota}$ is the direct limit of $(M_{\alpha},\sigma_{\alpha,\beta}:\alpha<\beta<\iota)$, $\sigma_{\alpha,\iota}:M_{\alpha}\rightarrow M_{\iota}$ is the direct limit embedding, and $\sigma_{\iota}:M_{\iota}\rightarrow U_{\chi}$ is the canonical realizability embedding given by $\sigma_{\iota}(x)=\sigma_{\alpha}(\bar{x})$ where $\alpha<\iota$ and $\bar{x}\in M_{\alpha}$ are such that $\sigma_{\alpha,\iota}(\bar{x})=x$. We call $M_{\iota}$ the direct limit of ${\sf{dm}}$.